OpenStudy (anonymous):

7+2lnX=2 solve for x?

OpenStudy (anonymous):

help please!!!

OpenStudy (anonymous):

\[e ^{-5/2}\]

OpenStudy (anonymous):

it is a decimal answer..... thats all i know

OpenStudy (lalaly):

\[2\ln x = -5\]\[\ln x=-5/2\]\[e ^{-5/2}\] = 0.820

OpenStudy (lalaly):

x= 0.820